# Maximal variation of linear systems

### v2

@@ Abstract.
Let $X$ be a smooth projective variety over $\mathbb{C}$, and $L$ a line bundle on $X$. We say that the linear system $\lvert{L}\rvert$ has maximal variation if its elements have the maximum number $\dim\lvert{L}\rvert$ of moduli. We discuss four cases where this situation is expected: hypersurfaces, double coverings of $\mathbb{P}^{n}$, K3 surfaces, and abelian varieties.
@@ Introduction
There is a simple differential criterion for this to happen, namely $H^{0}(T_{X|S})=0$ (§1). However this turns out to be surprisingly difficult to check. We will discuss four cases where maximal variation is expected: hypersurfaces (§2), double coverings of $\mathbb{P}^{n}$ (§3), K3 surfaces (§4), and abelian varieties (§5). We prove indeed that the variation is maximal in the latter case, while we have only partial results in the first three cases. Finally we discuss some more examples and counter-examples (§6 and 7).
@@ 3. Double coverings of $\mathbb{P}^{n}$
3) is proved in [C-D, Corollary 4.21], 4) in [D-H, Theorem 1.1].∎
@@ 6. Some more examples
We now observe that $X$ is a minimal surface: if $R$ is a smooth curve in $X$, $\Omega^{1}_{R}$ is a quotient of $\Omega^{1}_{X}$, hence it must be globally generated, which implies $g(R)\geq 1$. Browsing through the classification of surfaces we get the following cases:
@@ 7. Some counter-examples
$\bullet$ There are varieties of general type admitting a pencil with no moduli. Take for instance the Fermat hypersurface $F:\sum\limits_{i=0}^{n}x_{i}^{d}=0$. Consider the rational map $p:F\dasharrow\mathbb{P}^{1}$ given by $p(x_{0},\ldots,x_{n})=(x_{0},x_{1})$. Let $\hat{F}\rightarrow F$ be the blowing up of the codimension 2 subvariety given by $x_{0}=x_{1}=0$. Then $p$ extends to a morphism $\hat{p}:\hat{F}\rightarrow\mathbb{P}^{1}$ whose general fiber is isomorphic to the Fermat hypersurface of degree $d$ in $\mathbb{P}^{n-1}$. The corresponding linear system $\lvert{\hat{p}^{*}\mathcal{O}_{\mathbb{P}^{1}}(1)}\rvert$ has no moduli.
@@
It is probably overoptimistic to expect a positive answer, but it seems not obvious to find a counter-example.

### v3

@@ Abstract.
Let $X$ be a smooth projective variety over $\mathbb{C}$, and $L$ a line bundle on $X$. We say that the linear system $\lvert{L}\rvert$ has maximal variation if its elements have the maximum number $\dim\lvert{L}\rvert$ of moduli. We discuss some cases where this situation is expected: hypersurfaces, double coverings of $\mathbb{P}^{n}$, K3 and hyperkähler manifolds, and abelian varieties.
@@ Introduction
There is a simple differential criterion for this to happen, namely $H^{0}(T_{X|S})=0$ (§1). However this turns out to be surprisingly difficult to check. We will discuss some cases where maximal variation is expected: hypersurfaces (§2), double coverings of $\mathbb{P}^{n}$ (§3), K3 surfaces and hyperkähler manifolds (§4), and abelian varieties (§5). We prove indeed that the variation is maximal in the latter case, while we have only partial results in the former cases. Finally we discuss some more examples and counter-examples (§6 and 7).
@@ 4. K3 and hyperkähler
3) follows from [D-H, Corollary 5.3]; 4) is proved in [D-H, Theorem 1.1].∎
The case of higher-dimensional hyperkähler manifolds turns out to be much easier:
Proposition 6.
Let $X$ be a hyperkähler manifold of dimension >2>2. For any ample line bundle $L$ on $X$, the linear system $\lvert{L}\rvert$ has maximal variation.
Proof : Let $S\in\lvert{L}\rvert$. The exact sequence $0\rightarrow T_{X}\otimes L^{-1}\rightarrow T_{X}\rightarrow T_{X|S}\rightarrow 0$ gives rise to a cohomology exact sequence $H^{0}(T_{X})\rightarrow H^{0}(T_{X|S})\rightarrow H^{1}(T_{X}\otimes L^{-1})$. Since $T_{X}\cong\Omega^{1}_{X}$, we have $H^{0}(T_{X})=0$ and $H^{1}(T_{X}\otimes L^{-1})=0$ by the Akizuki-Nakano vanishing theorem, hence $H^{0}(T_{X|S})=0$.∎
@@ 6. Some more examples
We observe that $X$ is a minimal surface: if $R$ is a smooth curve in $X$, $\Omega^{1}_{R}$ is a quotient of $\Omega^{1}_{X}$, hence it must be globally generated, which implies $g(R)\geq 1$. Browsing through the classification of surfaces we get the following cases:
@@ 7. Some counter-examples
$\bullet$ There are varieties of general type admitting a linear system with no base point and no moduli. Take for instance the Fermat hypersurface $F:\sum\limits_{i=0}^{n}x_{i}^{d}=0$ in $\mathbb{P}^{n}$. For $1\leq r\leq n-2$, consider the rational map $p:F\dasharrow\mathbb{P}^{r}$ given by $p(x_{0},\ldots,x_{n})=\allowbreak(x_{0},x_{1},\ldots,x_{r})$. Let $\hat{F}\rightarrow F$ be the blowing up of the codimension $(r+1)$ subvariety of $F$ given by $x_{0}=\ldots=x_{r}=0$. Then $p$ extends to a morphism $\hat{p}:\hat{F}\rightarrow\mathbb{P}^{r}$ whose general fiber is isomorphic to the Fermat hypersurface of degree $d$ in $\mathbb{P}^{n-r}$. The corresponding linear system $\lvert{\hat{p}^{*}\mathcal{O}_{\mathbb{P}^{1}}(1)}\rvert$ has dimension $r$, and no moduli.
@@
It is perhaps overoptimistic to expect a positive answer, but it seems not obvious to find a counter-example.